OpenStudy (anonymous):

what is the coefficant x^2 y^3 in the expansion of (2x+y)^5 A.2 B.5 C.40 D.80 C.it does not exist

OpenStudy (cwrw238):

this is a Binomial expansion

OpenStudy (anonymous):

yea that is that is what the quiz is on can u help me

OpenStudy (cwrw238):

sorry i got that wrong its (a = b)^n = a^n + nC1 a(n-1) b^1 + nC2 a(n-2) b ^2 + ... in your case n = 5 a = 2x and b = y

OpenStudy (cwrw238):

oohh thats a^(n-1) and a^(n-2)

OpenStudy (anonymous):

so whats the answer ?

OpenStudy (cwrw238):

so for example your second term is 5C1 (2x)^(5-1) (y)^1 so coefficient of 2nd term is 5C1* 2^4 = 5 * 16 = 80

OpenStudy (anonymous):

ok i got another ?

OpenStudy (cwrw238):

can you do this one now - you need term which includes x2 y^3

OpenStudy (anonymous):

can u help me get the answers iv only go 10 more minutes on this quiz then we can work on them

OpenStudy (anonymous):

The bionomial expansion of (2x-y)^3 is A.x^3-6x^2y+12y^2-8y^3 B.x^3-3x^2y+3xy^2-y^3 C.2^3(X^3-3X^2Y+3XY^2-Y^3) D.8x^3-12x^2y+6x^2-y^3

OpenStudy (cwrw238):

if we write the term as 5Cr (2x)^(n-r) y ^r we have r = 3 because we need x^2 so your coefficient is 5C3 * 2^2 = 10 * 2^2 * 1 = 40

OpenStudy (cwrw238):

(a + b)^3 = a^3 + 3ab^2 + 3a^2b + b^3 replace a by 2x and b by -y and see which one you get

OpenStudy (anonymous):

so what would the answer be